OpenStudy (konradzuse):

The time to complete a standardized exam is approximately Normal with a mean of 70 minutes and a standard deviation of 10 minutes. Using the 68–95–99.7 rule, if students are given 90 minutes to complete the exam, what percent of students will not finish? A. 32% B. 5% C. 2.5% D. 0.0015%

OpenStudy (konradzuse):

I say C

OpenStudy (konradzuse):

90 mins is 2 deviations which leaves 2.5 %

OpenStudy (anonymous):

Lol you're answering your own questions. Too op.

OpenStudy (konradzuse):

Like I said I'm bjust double checking myself if I'm wrong then it's easier tocorrect it :).

OpenStudy (konradzuse):

Was I correct on this one? C?

OpenStudy (anonymous):

Lol i'm not even in standard devs yet. Only 17 lolz.

OpenStudy (konradzuse):

oh okay :)

OpenStudy (konradzuse):

I did this with calc 2, when byou try to answer it on your own it helps sink in.

OpenStudy (konradzuse):

@jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

99.7% of the students will have times from 70-2*10 = 50 min to 70+2*10 = 90 min (100 - 99.7)/2 = 0.15% will have times below 50 min So 99.7+0.15 = 99.85% will have times below 90 min This means that 100 - 99.85 = 0.15% of the students will not be able to finish the exam.

OpenStudy (konradzuse):

:O

OpenStudy (konradzuse):

however that isn't a choice.... 0.0015% is... LOL.

OpenStudy (konradzuse):

>(

OpenStudy (anonymous):

I like the process... is 95 supposed to correspond to 2 standard deviations?

OpenStudy (konradzuse):

mhm

OpenStudy (konradzuse):

I keep getting my formulas konfused...

OpenStudy (anonymous):

then it would be C using his process, but replacing 99.7 with 95.

jimthompson5910 (jim_thompson5910):

oh right...lol my bad

jimthompson5910 (jim_thompson5910):

99.7 is 3 std deviations away

jimthompson5910 (jim_thompson5910):

95 is 2

OpenStudy (konradzuse):

I basically thoght 10 is the dev.

jimthompson5910 (jim_thompson5910):

10 is the std dev

OpenStudy (konradzuse):

so then it's m + 2sigma

OpenStudy (konradzuse):

so 70+20 = 90 and that leaves 2.35 and .15

OpenStudy (konradzuse):

which is 2.5%

OpenStudy (konradzuse):

So I was right >( no medals >(

OpenStudy (anonymous):

;)